# Nonequispaced fast Fourier transforms for bandlimited functions

### v1

@@ 1 Introduction
with given Fourier coefficients $\hat{f}_{\boldsymbol{k}}\in\mathbb{C}$, $\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}$, at given nonequispaced points $\boldsymbol{x}_{j}\in\mathbb{T}^{d}$, $j=1,\dots,N$, $N\in\mathbb{N}$, where for $\boldsymbol{M}\coloneqq(M,\dots,M)^{T}$, $M\in 2\mathbb{N}$, we define the index set $\mathcal{I}_{\boldsymbol{M}}\coloneqq\mathbb{Z}^{d}\cap\left[-\tfrac{M}{2},\tfrac{M}{2}\right)^{d}$ with cardinality $|\mathcal{I}_{\boldsymbol{M}}|=M^{d}$, and $\mathbb{T}^{d}\coloneqq\mathbb{R}^{d}\setminus\mathbb{Z}^{d}$ denotes the $d$-dimensional torus with $d\in\mathbb{N}$.
@@
from given measurements $\hat{f}(\boldsymbol{k})\in\mathbb{C}$, $\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}$, of its functions Fourier transform $\hat{f}$.
@@ 2 The NFFT
where the inner product of two vectors shall be defined as usual as $\boldsymbol{k}\boldsymbol{x}\coloneqq k_{1}x_{1}+\dots+k_{d}x_{d}$. A fast approximate algorithm, the so-called nonequispaced fast Fourier transform (NFFT), can be summarized as follows, see e. g. [6, 2, 27, 9, 12] or [21, pp. 413–417].
@@
For $d,N\in\mathbb{N}$ and $\boldsymbol{M}=(M_{1},\dots,M_{d})^{\top}\in(2\mathbb{N})^{d}$ let $\boldsymbol{x}_{j}\in\mathbb{T}^{d},\,j=1,\dots,N,$ be given nodes as well as $\hat{f}_{\boldsymbol{k}}\in\mathbb{C}$, $\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}$, given Fourier coefficients. Furthermore, we are given the $d$-dimensional oversampling factor $\boldsymbol{\sigma}\geq\boldsymbol{1}_{d}$, the vector $\boldsymbol{M}_{\boldsymbol{\sigma}}\coloneqq((M_{\sigma})_{1},\dots,(M_{\sigma})_{d})^{\top}$ with $2\mathbb{N}\ni(M_{\sigma})_{t}\coloneqq 2\lceil\,\lceil\sigma_{t}M_{t}\rceil/2\,\rceil$, as well as the window function $\varphi$, the truncated function $\varphi_{m}$ with $m\ll(M_{\sigma})_{t}$, $t=1,\dots,d$, and their $1$-periodic versions $\tilde{\varphi}$ and $\tilde{\varphi}_{m}$.
@@
      Compute the nonzero values $\tilde{\varphi}_{m}\left(\boldsymbol{x}_{j}-\boldsymbol{M}_{\boldsymbol{\sigma}}^{-1}\odot\boldsymbol{\ell}\right)$ for $j=1,\dots,N,$ as well as $\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}},m}(\boldsymbol{x}_{j})$, cf. (2.5).
@@
    Compute
    $g_{\boldsymbol{\ell}}\coloneqq\frac{1}{|\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}}|}\sum_{\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}}\hat{g}_{\boldsymbol{k}}\,\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}(\boldsymbol{M}_{\boldsymbol{\sigma}}^{-1}\odot\,\boldsymbol{\ell})},\quad\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}},$
    by means of a $d$-variate iFFT. $\mathcal{O}(|\mathcal{I}_{\boldsymbol{M}}|\log(|\mathcal{I}_{\boldsymbol{M}}|))$
@@
    $\tilde{f}_{j}\coloneqq\sum_{\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}},m}(\boldsymbol{x}_{j})}g_{\boldsymbol{\ell}}\,\tilde{\varphi}_{m}\left(\boldsymbol{x}_{j}-\boldsymbol{M}_{\boldsymbol{\sigma}}^{-1}\odot\boldsymbol{\ell}\right),\quad j=1,\dots,N.$
  Output: $\tilde{f}_{j}\approx f(\boldsymbol{x}_{j})$, $j=1,\dots,N$, cf. (2.1). Complexity: $\mathcal{O}(|\mathcal{I}_{\boldsymbol{M}}|\log(|\mathcal{I}_{\boldsymbol{M}}|)+N)$
@@
$\boldsymbol{F}\coloneqq\left(\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}(\boldsymbol{M}_{\boldsymbol{\sigma}}^{-1}\odot\,\boldsymbol{\ell})}\right)_{\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}},\,\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}}\ \in\mathbb{C}^{|\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}}|\times|\mathcal{I}_{\boldsymbol{M}}|},$ (2.3)
and the sparse matrix
$\boldsymbol{B}\coloneqq\bigg{(}\tilde{\varphi}_{m}\left(\boldsymbol{x}_{j}-\boldsymbol{M}_{\boldsymbol{\sigma}}^{-1}\odot\boldsymbol{\ell}\right)\bigg{)}_{j=1,\,\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}}}^{N}\ \in\mathbb{R}^{N\times|\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}}|},$ (2.4)
@@
$\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}},m}(\boldsymbol{x}_{j})\coloneqq\left\{\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}}\colon\exists\boldsymbol{z}\in\mathbb{Z}^{d}\text{ with }-m\cdot\boldsymbol{1}_{d}\leq\boldsymbol{M}_{\boldsymbol{\sigma}}\odot(\boldsymbol{x}_{j}+\boldsymbol{z})-\boldsymbol{\ell}\leq m\cdot\boldsymbol{1}_{d}\right\}$ (2.5)
each row of $\boldsymbol{B}$ contains at most $(2m+1)^{d}$ nonzeros, the NFFT in Algorithm 2.1 can be formulated in matrix-vector notation as $\boldsymbol{A}\approx\boldsymbol{B}\boldsymbol{F}\boldsymbol{D}$, cf. [21, p. 419]. This is to say, using the definition of the matrices, the NFFT performs the approximation
$\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}\boldsymbol{x}_{j}}\approx\frac{1}{|\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}}|\cdot\hat{\varphi}(\boldsymbol{k})}\sum_{\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}},m}(\boldsymbol{x}_{j})}\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}(\boldsymbol{M}_{\boldsymbol{\sigma}}^{-1}\odot\,\boldsymbol{\ell})}\,\tilde{\varphi}_{m}\left(\boldsymbol{x}_{j}-\boldsymbol{M}_{\boldsymbol{\sigma}}^{-1}\odot\boldsymbol{\ell}\right)$ (2.6)
@@ 3 Regularized Shannon sampling formulas
is contained in $\left[-\frac{M}{2},\frac{M}{2}\right]^{d}$. The space of all bandlimited functions with shall be denoted by
@@
${\mathcal{B}}_{M/2}(\mathbb{R}^{d})\subseteq L_{2}(\mathbb{R}^{d})\cap C_{0}(\mathbb{R}^{d})\cap C^{\infty}(\mathbb{R}^{d}).$ (3.2)
cf. [13, Lemma 4.1]. Thus, the Fourier inversion theorem, see e. g. [21, Theorem 2.23], guarantees that the inverse Fourier transform of $f$ can be written as given in (1.1).
@@
where the $\mathrm{sinc}$ function is given by
$\mathrm{sinc}(\boldsymbol{x})\coloneqq\prod_{t=1}^{d}\mathrm{sinc}(x_{t}),\quad\mathrm{sinc}(x)\coloneqq\left\{\begin{array}[]{ll}\frac{\sin x}{x}&\colon x\in\mathbb{R}\setminus\{0\},\\[4.30554pt] 1&\colon x=0,\end{array}\right.$
@@
Unfortunately, the numerical use of this classical Whittaker–Kotelnikov–Shannon sampling series (3.3) is limited, since it requires infinitely many samples, which is impossible in practice, and its truncated version is not a good approximation due to the slow decay of the $\mathrm{sinc}$ function, see [10]. In addition to this rather poor convergence, it is known, see [7, 8, 5], that in the presence of noise in the samples $f(\boldsymbol{k})$, $\boldsymbol{k}\in\mathbb{Z}^{d}$, the convergence of Shannon sampling series (3.3) may even break down completely.
@@
$\varphi_{\sinh}(x)\coloneqq\begin{cases}\frac{1}{\sinh\beta}\,\sinh\Big{(}\beta\,\sqrt{1-\big{(}\tfrac{Lx}{m}\big{)}^{2}}\,\Big{)}&\quad\colon x\in\left[-\tfrac{m}{L},\,\tfrac{m}{L}\right],\\ 0&\quad\colon x\in\mathbb{R}\setminus\left[-\tfrac{m}{L},\,\tfrac{m}{L}\right],\end{cases}$ (3.4)
@@
$\varphi_{\mathrm{cKB}}(x)\coloneqq\begin{cases}\frac{1}{I_{0}(\beta)-1}\,\bigg{(}I_{0}\Big{(}\beta\,\sqrt{1-\big{(}\tfrac{Lx}{m}\big{)}^{2}}\,\Big{)}-1\bigg{)}&\quad\colon x\in\left[-\tfrac{m}{L},\,\tfrac{m}{L}\right],\\ 0&\quad\colon x\in\mathbb{R}\setminus\left[-\tfrac{m}{L},\,\tfrac{m}{L}\right],\end{cases}$
with certain $\beta>0$, where $I_{0}$ denotes the modified Bessel function of first kind. Note that these window functions are well-studied, in the context of the NFFT, see e. g. [22].
@@
since by assumption $\varphi(0)=1$ and $\mathrm{sinc}(\pi(k-\ell))=\delta_{k,\ell}$ for all $k,\,\ell\in\mathbb{Z}$ with the Kronecker symbol $\delta_{k,\ell}$. Further we assume that the samples $f\big{(}\frac{\boldsymbol{\ell}}{L}\big{)}$, $\boldsymbol{\ell}\in\mathbb{Z}^{d}$, fulfill the condition
$\sum_{\boldsymbol{\ell}\in\mathbb{Z}^{d}}\big{|}f\big{(}\tfrac{\boldsymbol{\ell}}{L}\big{)}\big{|}<\infty.$
Then it is known that the regularized Shannon sampling formula $R_{\varphi,m}f$ in (3.5) with suitable window function $\varphi\in\Phi_{m,L}$ yields a good approximation of $f$, cf. [14, 15, 13].
@@ 4 NFFT-like procedure for bandlimited functions
Now assume we are given the values $\hat{f}(\boldsymbol{k})$, $\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}$, of the Fourier transform (1.1) of a bandlimited function $f\in\mathcal{B}_{M/2}(\mathbb{R}^{d})$, and we are looking for function evaluations $f(\boldsymbol{x}_{j})$ at given nonequispaced points $\boldsymbol{x}_{j}$, $j=1,\dots N$.
In order to compute these values, we make use of the regularized Shannon sampling formulas, see Section 3. Inserting the approximation (3.5) into the Fourier transform (3.1) and using the definition of the regularized $\mathrm{sinc}$ function $\psi(\boldsymbol{x})\coloneqq\mathrm{sinc}(L\pi\boldsymbol{x})\,\varphi(\boldsymbol{x})$, we have
@@
$=\bigg{(}\sum_{\boldsymbol{\ell}\in\mathbb{Z}^{d}}f\big{(}\tfrac{\boldsymbol{\ell}}{L}\big{)}\,\mathrm{e}^{-2\pi\mathrm{i}\boldsymbol{v}\boldsymbol{\ell}/L}\bigg{)}\cdot\hat{\psi}(\boldsymbol{v}).$ (4.1)
@@
To finally approximate the samples $f(\boldsymbol{x}_{j})$, $j=1,\dots,N$, we make use of the regularized Shannon sampling formula (3.5). Note that since we assumed that the window function $\varphi\in\Phi_{m,L}$ is compactly supported, the computation of $(R_{\varphi,m}f)(\boldsymbol{x})$ for fixed $\boldsymbol{x}\in\mathbb{R}^{d}\setminus\tfrac{1}{L}\,\mathbb{Z}^{d}$ requires only $(2m+1)^{d}$ samples $f\big{(}\tfrac{\boldsymbol{\ell}}{L}\big{)}$. However, we have already encountered that (4.5) can only be used to approximate $f\big{(}\tfrac{\boldsymbol{\ell}}{L}\big{)}$ for $\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{L}}$ in order to avoid aliasing in the computation of the inverse Fourier transform in (4.5). Thereby, we are confronted with a limitation of the feasible interval for the points $\boldsymbol{x}_{j}\in[-\tfrac{1}{2}+\tfrac{m}{L},\tfrac{1}{2}-\tfrac{m}{L})^{d}$, $j=1,\dots,N$, since only in this case exclusively the evaluations $f\big{(}\tfrac{\boldsymbol{\ell}}{L}\big{)}$, $\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{L}}$, are needed for the computation. Hence, the final approximation is computed by
$(R_{\varphi,m}f)(\boldsymbol{x}_{j})\approx f_{j}$ $\coloneqq\sum_{\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{L}}}\vartheta_{\boldsymbol{\ell}}\,\psi\big{(}\boldsymbol{x}_{j}-\tfrac{\boldsymbol{\ell}}{L}\big{)}$
$=\sum_{\boldsymbol{\ell}\in\mathcal{J}_{\boldsymbol{L},m}(\boldsymbol{x}_{j})}\vartheta_{\boldsymbol{\ell}}\,\psi\big{(}\boldsymbol{x}_{j}-\tfrac{\boldsymbol{\ell}}{L}\big{)},$
@@
contains at most $(2m+1)^{d}$ entries for each fixed $\boldsymbol{x}_{j}$, cf. (2.5). Thus, the obtained algorithm can be summarized as follows, cf. [13, Algorithm 5.16].
@@
For $d,m,N\in\mathbb{N}$ and $M\in 2\mathbb{N}$ let $\boldsymbol{x}_{j}\in[-\tfrac{1}{2}+\tfrac{m}{L},\tfrac{1}{2}-\tfrac{m}{L})^{d}$, $j=1,\dots,N$, be given nodes as well as $\hat{f}(\boldsymbol{k})\in\mathbb{C}$, $\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}$, given evaluations of the Fourier transform of the bandlimited function $f\in\mathcal{B}_{M/2}(\mathbb{R}^{d})$. Furthermore, we are given the oversampling parameter $\lambda\geq 0$ with $L=M(1+\lambda)\in\mathbb{N}$, as well as the window function $\varphi\in\Phi_{m,L}$, the corresponding regularized $\mathrm{sinc}$ function $\psi$ and its Fourier transform $\hat{\psi}$.
@@
    Compute
@@
    by means of a $d$-variate iFFT. $\mathcal{O}(|\mathcal{I}_{\boldsymbol{M}}|\log(|\mathcal{I}_{\boldsymbol{M}}|))$
@@
  Output: $f_{j}\approx f(\boldsymbol{x}_{j})$, cf. (1.1), $j=1,\dots,N$. Complexity: $\mathcal{O}(|\mathcal{I}_{\boldsymbol{M}}|\log(|\mathcal{I}_{\boldsymbol{M}}|)+N)$

### v2

@@ 1 Introduction
with given Fourier coefficients $\hat{f}_{\boldsymbol{k}}\in\mathbb{C}$, $\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}$, at given nonequispaced points $\boldsymbol{x}_{j}\in\mathbb{T}^{d}$, $j=1,\dots,N$, $N\in\mathbb{N}$, where for $M\in 2\mathbb{N}$ we define the index set $\mathcal{I}_{\boldsymbol{M}}\coloneqq\mathbb{Z}^{d}\cap\left[-\tfrac{M}{2},\tfrac{M}{2}\right)^{d}$ with cardinality $|\mathcal{I}_{\boldsymbol{M}}|=M^{d}$, and $\mathbb{T}^{d}\coloneqq\mathbb{R}^{d}/\mathbb{Z}^{d}$, $d\in\mathbb{N}$, denotes the $d$-dimensional torus.
@@
from given measurements $\hat{f}(\boldsymbol{k})\in\mathbb{C}$, $\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}$, of its Fourier transform $\hat{f}$.
@@ 2 The NFFT
where the inner product shall be defined as usual as $\boldsymbol{k}\boldsymbol{x}\coloneqq k_{1}x_{1}+\dots+k_{d}x_{d}$. A fast approximate algorithm, the so-called nonequispaced fast Fourier transform (NFFT), can be summarized as follows, see e. g. [6, 2, 29, 10, 13] or [22, pp. 413–417].
@@
For $d,N\in\mathbb{N}$ and $M\in 2\mathbb{N}$ let $\boldsymbol{x}_{j}\in\mathbb{T}^{d},\,j=1,\dots,N,$ be given nodes as well as $\hat{f}_{\boldsymbol{k}}\in\mathbb{C}$, $\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}$, given Fourier coefficients. Furthermore, we are given the oversampling factor $\sigma\geq 1$ with $M_{\sigma}\coloneqq 2\lceil\,\lceil\sigma M\rceil/2\,\rceil\in 2\mathbb{N}$, as well as the window function $\varphi$, the truncated function $\varphi_{m}$ with $m\ll M_{\sigma}$, and their $1$-periodic versions $\tilde{\varphi}$ and $\tilde{\varphi}_{m}$.
@@
      Compute the values $\tilde{\varphi}_{m}\big{(}\boldsymbol{x}_{j}-\frac{\boldsymbol{\ell}}{M_{\boldsymbol{\sigma}}}\big{)}$ for $j=1,\dots,N,$ as well as $\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{M_{\sigma}},m}(\boldsymbol{x}_{j})$, cf. (2.5).
@@
    Compute $\mathcal{O}(|\mathcal{I}_{\boldsymbol{M}}|\log(|\mathcal{I}_{\boldsymbol{M}}|))$
    $g_{\boldsymbol{\ell}}\coloneqq\frac{1}{|\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}}|}\sum_{\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}}\hat{g}_{\boldsymbol{k}}\,\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}\boldsymbol{\ell}/M_{\sigma}},\quad\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}},$
    by means of a $d$-variate iFFT.
@@
    $\tilde{f}_{j}\coloneqq\sum_{\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}},m}(\boldsymbol{x}_{j})}g_{\boldsymbol{\ell}}\,\tilde{\varphi}_{m}\big{(}\boldsymbol{x}_{j}-\tfrac{\boldsymbol{\ell}}{M_{\sigma}}\big{)},\quad j=1,\dots,N.$
  Output: $\tilde{f}_{j}\approx f(\boldsymbol{x}_{j})$ Complexity: $\mathcal{O}(|\mathcal{I}_{\boldsymbol{M}}|\log(|\mathcal{I}_{\boldsymbol{M}}|)+N)$
@@
$\boldsymbol{F}\coloneqq\left(\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}\boldsymbol{\ell}/M_{\boldsymbol{\sigma}}}\right)_{\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}},\,\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}}\in\mathbb{C}^{|\mathcal{I}_{\boldsymbol{M_{\sigma}}}|\times|\mathcal{I}_{\boldsymbol{M}}|},$ (2.3)
and the $(2m+1)^{d}$-sparse matrix
$\boldsymbol{B}\coloneqq\bigg{(}\tilde{\varphi}_{m}\big{(}\boldsymbol{x}_{j}-\tfrac{\boldsymbol{\ell}}{M_{\sigma}}\big{)}\bigg{)}_{j=1,\,\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}}}^{N}\in\mathbb{R}^{N\times|\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}}|},$ (2.4)
@@
$\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}},m}(\boldsymbol{x}_{j})$$\coloneqq\left\{\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}}\colon\exists\boldsymbol{z}\in\mathbb{Z}^{d}\text{ with }\right.\left.-m\cdot\boldsymbol{1}_{d}\leq M_{\sigma}\cdot(\boldsymbol{x}_{j}+\boldsymbol{z})-\boldsymbol{\ell}\leq m\cdot\boldsymbol{1}_{d}\right\}$ (2.5)
each row of $\boldsymbol{B}$ contains at most $(2m+1)^{d}$ nonzeros, the NFFT in Algorithm 2.1 can be formulated in matrix-vector notation as $\boldsymbol{A}\approx\boldsymbol{B}\boldsymbol{F}\boldsymbol{D}$, cf. [22, p. 419]. This is to say, using the definition of these matrices, the NFFT performs the approximation
$\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}\boldsymbol{x}_{j}}\approx\sum_{\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}},m}(\boldsymbol{x}_{j})}\frac{\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}\boldsymbol{\ell}/M_{\sigma}}\,\tilde{\varphi}_{m}\big{(}\boldsymbol{x}_{j}-\tfrac{\boldsymbol{\ell}}{M_{\sigma}}\big{)}}{|\mathcal{I}_{\boldsymbol{M_{\boldsymbol{\sigma}}}}|\cdot\hat{\varphi}(\boldsymbol{k})}$ (2.6)
@@ 3 Regularized Shannon sampling formulas
is contained in $\left[-\frac{M}{2},\frac{M}{2}\right]^{d}$. The space of all bandlimited functions with bandwidth $M$ shall be denoted by
@@
${\mathcal{B}}_{M/2}(\mathbb{R}^{d})\subseteq L_{2}(\mathbb{R}^{d})\cap C_{0}(\mathbb{R}^{d})\cap C^{\infty}(\mathbb{R}^{d}),$ (3.2)
cf. [14, Lemma 4.1]. Thus, the Fourier inversion theorem, see e. g. [22, Theorem 2.23], guarantees that the inverse Fourier transform of $f$ can be written as given in (1.1).
@@
where the $\mathrm{sinc}$ function is given by $\mathrm{sinc}(\boldsymbol{x})\coloneqq\prod_{t=1}^{d}\mathrm{sinc}(x_{t})$ with
$\mathrm{sinc}(x)\coloneqq\left\{\begin{array}[]{ll}\frac{\sin x}{x}&\colon x\in\mathbb{R}\setminus\{0\},\\[4.30554pt] 1&\colon x=0.\end{array}\right.$
@@
Unfortunately, the numerical use of this classical Whittaker–Kotelnikov–Shannon sampling series (3.3) is limited, since it requires infinitely many samples, which is impossible in practice, and its truncated version is not a good approximation due to the slow decay of the $\mathrm{sinc}$ function, see [11]. In addition to this rather poor convergence, it is known, see [8, 9, 5], that in the presence of noise in the samples $f\big{(}\tfrac{\boldsymbol{\ell}}{L}\big{)}$, $\boldsymbol{\ell}\in\mathbb{Z}^{d}$, the convergence of the Shannon sampling series (3.3) may even break down completely.
@@
$\!\!\varphi_{\sinh}(x)\coloneqq\frac{1}{\sinh\beta}\,\sinh\!\Big{(}\beta\sqrt{1-\big{(}\tfrac{Lx}{m}\big{)}^{2}}\,\Big{)}\,\chi_{\left[-\tfrac{m}{L},\,\tfrac{m}{L}\right]}(x)\!$ (3.4)
@@
$\varphi_{\mathrm{cKB}}(x)\coloneqq\frac{\bigg{(}I_{0}\Big{(}\beta\,\sqrt{1-\big{(}\tfrac{Lx}{m}\big{)}^{2}}\,\Big{)}-1\bigg{)}}{I_{0}(\beta)-1}\,\,\chi_{\left[-\tfrac{m}{L},\,\tfrac{m}{L}\right]}(x)$
with certain $\beta>0$, where $I_{0}$ denotes the modified Bessel function of first kind. Note that these window functions are well-studied in the context of the NFFT, see e. g. [23].
@@
since by assumption $\varphi(0)=1$ and $\mathrm{sinc}(\pi(k-\ell))=\delta_{k,\ell}$ for all $k,\,\ell\in\mathbb{Z}$ with the Kronecker symbol $\delta_{k,\ell}$. Then it is known that the regularized Shannon sampling formula $R_{\varphi,m}f$ in (3.5) with suitable window function $\varphi\in\Phi_{m,L}$ yields a good approximation of $f$, cf. [15, 16, 14].
@@ 4 NFFT-like procedure for bandlimited functions
Now assume we are given the values $\hat{f}(\boldsymbol{k})$, $\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}$, of the Fourier transform (1.1) of a bandlimited function $f\in\mathcal{B}_{M/2}(\mathbb{R}^{d})$, and we are looking for function evaluations $f(\boldsymbol{x}_{j})$ at given nonequispaced points $\boldsymbol{x}_{j}$, $j=1,\dots N$. Further we assume that the function $f$ fulfills the condition
$\sum_{\boldsymbol{\ell}\in\mathbb{Z}^{d}}\big{|}f\big{(}\tfrac{\boldsymbol{\ell}}{L}\big{)}\big{|}<\infty,$ (4.1)
such that the Fourier series
$\hat{f}(\boldsymbol{v})=\frac{1}{L^{d}}\sum_{\boldsymbol{\ell}\in\mathbb{Z}^{d}}f\big{(}\tfrac{\boldsymbol{\ell}}{L}\big{)}\,\mathrm{e}^{-2\pi\mathrm{i}\boldsymbol{\ell}\boldsymbol{v}/L},\quad\boldsymbol{v}\in\left[-\tfrac{L}{2},\,\tfrac{L}{2}\right]^{d},$
converges absolutely and uniformly, see [16, p. 12], and thus the considered problem is well-defined. Note that by [27, Lemma 2] it is known that $f\in L_{1}(\mathbb{R}^{d})$ fulfill the condition (4.1). Moreover, $f\in L_{1}(\mathbb{R}^{d})$ directly implies $\hat{f}\in C_{0}(\mathbb{R}^{d})$.
Remark 4.1.
Note that the recent work [7] derived error estimates for a familiar problem, where however for functions $f\in C(\mathbb{R})$ satisfying certain decay and smoothness conditions and equispaced points $\boldsymbol{x}_{j}=\frac{2j-N}{2N}$, $j=1,\dots,N$, simply the FFT can be used.
In order to compute the values $f(\boldsymbol{x}_{j})$, $j=1,\dots N$, we aim to make use of the regularized Shannon sampling formulas, see Section 3. Inserting the approximation (3.5) into the Fourier transform (3.1) and using the definition of the regularized $\mathrm{sinc}$ function
$\psi(\boldsymbol{x})\coloneqq\mathrm{sinc}(L\pi\boldsymbol{x})\,\varphi(\boldsymbol{x}),$ (4.2)
we have
@@
$=\bigg{(}\sum_{\boldsymbol{\ell}\in\mathbb{Z}^{d}}f\big{(}\tfrac{\boldsymbol{\ell}}{L}\big{)}\,\mathrm{e}^{-2\pi\mathrm{i}\boldsymbol{v}\boldsymbol{\ell}/L}\bigg{)}\cdot\hat{\psi}(\boldsymbol{v}),$ (4.3)
@@
To finally approximate the samples $f(\boldsymbol{x}_{j})$, $j=1,\dots,N$, we make use of the regularized Shannon sampling formula (3.5). Note that since we assumed that the window function $\varphi\in\Phi_{m,L}$ is compactly supported, the computation of $(R_{\varphi,m}f)(\boldsymbol{x})$ for fixed $\boldsymbol{x}\in\mathbb{R}^{d}\setminus\tfrac{1}{L}\,\mathbb{Z}^{d}$ requires only $(2m+1)^{d}$ samples $f\big{(}\tfrac{\boldsymbol{\ell}}{L}\big{)}$. However, we have already encountered that (4.7) can only be used to approximate $f\big{(}\tfrac{\boldsymbol{\ell}}{L}\big{)}$ for $\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{L}}$ in order to avoid aliasing in the computation of the inverse Fourier transform in (4.7). Thereby, we are confronted with a limitation of the feasible points to $\boldsymbol{x}_{j}\in[-\tfrac{1}{2}+\tfrac{m}{L},\tfrac{1}{2}-\tfrac{m}{L})^{d}$, $j=1,\dots,N$, since only in this case exclusively the evaluations $f\big{(}\tfrac{\boldsymbol{\ell}}{L}\big{)}$, $\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{L}}$, are needed for the computation. Hence, the final approximation is computed by
$(R_{\varphi,m}f)(\boldsymbol{x}_{j})\approx f_{j}$ $\coloneqq\sum_{\boldsymbol{\ell}\in\mathcal{I}_{\boldsymbol{L}}}\vartheta_{\boldsymbol{\ell}}\,\psi\big{(}\boldsymbol{x}_{j}-\tfrac{\boldsymbol{\ell}}{L}\big{)}=\sum_{\boldsymbol{\ell}\in\mathcal{J}_{\boldsymbol{L},m}(\boldsymbol{x}_{j})}\vartheta_{\boldsymbol{\ell}}\,\psi\big{(}\boldsymbol{x}_{j}-\tfrac{\boldsymbol{\ell}}{L}\big{)},$
@@
contains at most $(2m+1)^{d}$ entries for each fixed $\boldsymbol{x}_{j}$, cf. (2.5). Thus, the obtained algorithm can be summarized as follows, cf. [14, Algorithm 5.16].
@@
For $d,m,N\in\mathbb{N}$, $M\in 2\mathbb{N}$, and $L=M(1+\lambda)\in\mathbb{N}$ with oversampling parameter $\lambda\geq 0$ let $\boldsymbol{x}_{j}\in[-\tfrac{1}{2}+\tfrac{m}{L},\tfrac{1}{2}-\tfrac{m}{L})^{d}$, $j=1,\dots,N$, be given nodes as well as $\hat{f}(\boldsymbol{k})\in\mathbb{C}$, $\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}$, given evaluations of the Fourier transform of the bandlimited function $f\in\mathcal{B}_{M/2}(\mathbb{R}^{d})$. Furthermore, we are given the window function $\varphi\in\Phi_{m,L}$, the corresponding regularized $\mathrm{sinc}$ function $\psi$ in (4.2), and its Fourier transform $\hat{\psi}$.
@@
    Compute $\mathcal{O}(|\mathcal{I}_{\boldsymbol{M}}|\log(|\mathcal{I}_{\boldsymbol{M}}|))$
@@
    by means of a $d$-variate iFFT.
@@
  Output: $f_{j}\approx f(\boldsymbol{x}_{j})$ Complexity: $\mathcal{O}(|\mathcal{I}_{\boldsymbol{M}}|\log(|\mathcal{I}_{\boldsymbol{M}}|)+N)$